You improve your time running a course by 5 seconds in week one, by 3 seconds in week two, and by 4 seconds in week three. What is the mean weekly change in your running time?
1
Expert's answer
2019-04-18T03:02:58-0400

Mean weekly change: a=(5+3+4)/3=12/3=4 sec.